OpenStudy (waheguru):

Can someone help me solve this (x+2)^2 = 11^2+4^2 I want to solve it by expanding the (x+2)^2

hartnn (hartnn):

so, can you expand (x+2)^2 ?

OpenStudy (waheguru):

yes it becomes x^2+4x+4

hartnn (hartnn):

x^2+4x+4 = 11^2+4^2 whats 11^2+4^2 =... ?

OpenStudy (waheguru):

137

OpenStudy (waheguru):

x^2+4x = 133

OpenStudy (waheguru):

but what now from here ?

hartnn (hartnn):

x^2 +4x -133 =0 do you know how to solve quadratic equations ?

OpenStudy (waheguru):

no

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (waheguru):

Thanks,

hartnn (hartnn):

ok, welcome :) you can ask if you get stuck while getting the values of x